OpenStudy (anonymous):

how would i write the percentage of6/25

OpenStudy (anonymous):

6/25 x 4/4 = 24/100 = 24%

OpenStudy (anonymous):

Percent means something out of a 100 6/25 is like saying 6 out of 25. We can keep adding 6/25 until the denominator(25) = 100 or we can set it up as a proportion. We want to know what 6/25 is out of a 100. We can do this by using proportions or the method Rita used or by adding. You would set up the proportion like below 6/25 = x/100 Cross multiply 6 * 100 = 600 Divide 600 by 25 600/25 = 24% If we did this by adding we would set the problem up like so 6/25 + 6/25 + 6/25 + 6/25 = 24/100 24/100 means 24 out of 100 or 24%

OpenStudy (anonymous):

thanks XD